DarkChubstersYT:

Hello yall! I need help with math, Ive been posting math questions and nobody helps me exept My friends. Please help me!

Midnight97:

yeah i can help

DarkChubstersYT:

Arieonna:

@jhonyy9 can u help with math

DarkChubstersYT:

@jhonyy9 please

jhonyy9:

first x y 2×7=14 4×10=40 6×13=78 how you see it above how you think is or not is proportional x and y ? second x y 4×8=32 4×10=40 4×12=48 what is the difference between two cases ?

DarkChubstersYT:

Ohhhh TYSM!!